OpenStudy (anonymous):

How do i find the limit of (1/5^n)?

OpenStudy (anonymous):

hey

OpenStudy (rational):

plugin n = infinity

OpenStudy (anonymous):

yeah i forgot to put that

OpenStudy (rational):

1/5 = 0.2 1/5^2 = 0.04 ... 1/5^10 = 0.0000000000000000024 Notice that as we make "n" large, the value of 1/5^n is approaching 0. This should convince us that the limit of 1/5^n is 0.

OpenStudy (anonymous):